# Mediation, splitting and shellability in political structures modeled by simplicial complexes

### v1

@@ Abstract.
Modeling political structures by simplicial complexes, we investigate whether introducing a mediator into a substructure increases or decreases the stability of the overall structure. We prove theorems that quantify the stability of a political structure when $n$ mediators are introduced, either one by one or simultaneously. We also examine how the stability is affected when a single agent is split into two. In addition, stability is expressed in terms of the $h$-vector, and special attention is given to a class of political structures modeled by shellable simplicial complexes. In the latter context, we analyze weighted political structures and examples of political structures modeled by independence complexes of graphs.
@@
Acknowledgements. The authors would like to thank Ismar Volić for looking over the manuscript and providing helpful suggestions.
@@ 1. Introduction
In the specific case that is relevant to us here, the authors in [15] establish a natural correspondence (the map $\varphi$ following Definition 2.2) between the political structure (Definition 2.1) on the one hand and the simplicial complex (Definition 2.2) on the other.
@@
In this paper, we investigate the effects of introducing a mediator into a substructure of a political structure, motivated by an open question posed after [15, Proposition 3.15], as well as the effects of introducing $n$ mediators into a political structure. We also investigate what happens when instead of merging agents we have splitting of an agent - in other words, when a vertex in a simplicial complex is split into two new vertices. Such situations are natural to observe, bearing in mind, for example, the splitting of political parties into new parties. Furthermore, we describe the stability of a political structure modeled by a simplicial complex via the $h$-vector of the complex, with particular attention to the case of shellable complexes. In the end, we model political structures as independence complexes of graphs, and illustrate this with several examples.
@@ 2. Overview of some basic concepts and results
The cited literature - especially [1, 6, 15] - provides a detailed introduction to the terminology of political structures, as well as the necessary background on simplicial complexes. Accordingly, here we will limit ourselves to a brief review of key concepts and results that are directly relevant to the problems under consideration.
@@
Given a $d$-dimensional political structure $P$ with the set of agents $V=\{v_{0},v_{1},\ldots,v_{k}\}$,
@@
    where $\operatorname{st}(v_{i})$ is the star of $v_{i}$, i.e. the collection of all simplices of $P$ containing $a_{i}$, and $|\cdot|$ denotes the cardinality.
@@ 3. Effects of introducing a mediator into the substructure of a political structure
We made a simple observation that identifies a condition under which the stability of the structure increases or decreases, not restricted to the case $G\geq k+1$, thereby providing a more complete answer to the aforementioned question from [15]. We state it as a corollary to the previous proposition.
@@
Therefore, $G$ is a characteristic indicating how benevolent (acceptable) a mediator must be in order to increase stability. Like the main notion $\operatorname{stab}(P)$, it depends only on the total number of vertices and the number of simplices of $P$.
@@
where $S\in\{k+1,k+2,\ldots,2^{k+1}-1\}$. The function $g(S)$ is strictly decreasing. Consequently, as $S$ increases, the number of unaffected simplices $G<g(S)$ for which the introduction of a mediator leads to an increase in system stability decreases. In other words, the more ”developed” the structure – i.e., the greater the number of simplices or pre-existing compatibilities among agents – the larger the subcomplex (in terms of cardinality) into which the mediator must be introduced in order to achieve an increase in overall structure stability.
@@ 4. Effects of introducing many mediators into a political structure
As previously described, introducing a single mediator $v$ into a political structure $P$ yields the cone $CP=P*\{v\}$ on $P$. Introducing a mediator increases the stability of the structure, that is, $\operatorname{stab}(P)\leq\operatorname{stab}(CP)$. Introducing simultaneously two mediators $v$ and $w$ into $P$ yields the suspension $\Sigma P=P*\{v,w\}$, where $\{v,w\}$ is just the set of vertices $v$ and $w$. The authors in [15] claim that $\operatorname{stab}(P)\leq\operatorname{stab}(\Sigma P)$; however, this need not be true. Their argument rests on the fact that the suspension can be seen as the pushout of two copies of the cone $CP$ along $P$, together with the claim that a pushout increases stability. But the pushout does not in general increase stability. It does so only when no new vertices are introduced, i.e. when we work with $P_{1}\coprod_{S}P_{2}$, where $S$ is a subcomplex of both $P_{1}$ and $P_{2}$, and $P_{2}$ introduces no new vertices ($V(P_{2})\subseteq V(P_{1})$). In the case of the suspension, of course, new vertices are necessarily added. If $V(P_{1})\not\subseteq V(P_{2})$, one often encounters
@@
The number $\frac{k+2}{2^{k+1}}$ is small for large $k$. Therefore, introducing many mediators does not significantly increase the stability of the structure, especially in the case of structures with a larger number of agents.
@@ 5. Splitting and stability
Assume that $v$ is an agent (a vertex) of $P$ and that $v$ splits into two agents $a$ and $b$ (new vertices). Each of these new agents forms new coalitions, which can be described as the links $L_{a}$ and $L_{b}$ of the vertices $a$ and $b$, respectively. Therefore, the resulting political structure is
$P_{v}^{\prime}=(P\setminus\{v\})\amalg a*L_{a}\amalg b*L_{b}.$
@@
Therefore, $X$ is a characteristic described using the notions introduced in [15].
@@ 6. Shellable political structures and their stability
As an example the boundary of every simplicial polytope is shellable. From a topological perspective, a pure shellable simplicial complex of dimension $d$ is homotopy equivalent to a wedge of $d$-dimensional spheres or is contractible.
@@
If the political structure under consideration is shellable, the formula (4) becomes particularly useful. In that case, we can easily track how the stability changes by adding or deleting some facets (keeping the shellability). Adding (or removing) a new facet of type $i>1$ (there are no new vertices, i.e. agents), increases (or decreases) the corresponding entry $h_{i}$ by one. Consequently, the stability of $P$ increases (or decreases) by
@@
Consider what happens if we glue a facet that brings one new vertex, i.e., a facet of type $1$. Assume that $P=F_{1}\cup F_{2}\cup\cdots\cup F_{j}$ and let $F_{j+1}$ be a $d$-dimensional simplex of type 1, where $P$ has $k+1$ agents. If $\widetilde{S}$ is the total number of positive-dimensional simplices in $P$, then $\operatorname{stab}(P)=\frac{\widetilde{S}}{2^{k+1}-k-2}$. If the dimension of $P$ is $d$, then for
@@
If all complexes $P_{j}$ are shellable, and if $P_{j+1}$ is obtained from the previous one by adding a few new faces, we can easily calculate the stabilities and the weighted stability. That in itself is reasonable motivation to consider shellable structures in this context.
@@ 7. The independence complex of a graph as a political structure
Here we describe a natural way to model political structures using graphs. Consider a graph $G=(V,E)$ with $k+1$ vertices. Assume that the vertices of $G$ represent the agents. Let any two agents connected by an edge in $G$ be in conflict and therefore unable to appear together in any viable configuration. In this setting, the set of all viable configurations corresponds precisely to the independence complex $I(G)$ of the graph $G$, which is the simplicial complex whose faces are the independent sets of vertices of $G$. For a formal definition and more on independence complexes, including their connection to shellability, see [9].
@@
Now consider $k+1=2n$ agents labeled by the set $\{1,2,\ldots,n\}\cup\{-1,-2,\ldots,-n\}$. Each agent $i$ is in conflict solely with agent i-i, and these conflicts are represented by a graph with $n$ edges connecting $i$ and i-i for all $i\in[n]$. The corresponding political structure is $P=\partial C^{\triangle}_{n-1}$, the boundary of (n1)(n-1)-dimensional crosspolytope $C^{\triangle}_{n-1}$ (the dual of (n1)(n-1)-cube). The complex $\partial C^{\triangle}_{n-1}$ has $2^{n}$ facets, each corresponding to an ordered partition of $[n]$ into two subsets. For a partition $n]=A\sqcup B$, the facet associated to $(A,B)$ is defined by
@@ 8. Further questions
    According to the definition of stability of a political structure, two structures with the same number of vertices and the same number of $(\geq 1)$-dimensional simplices will have the same stability. However, these structures need not be topologically equivalent. Therefore, the stability defined in this manner is not a topological invariant, which makes it problematic if one aims to approach the problem from a more topological perspective.

### v2

# Stability of political structures modeled by simplicial complexes under mediation, splitting, and shellability

@@ Abstract.
Modeling political structures by simplicial complexes, we investigate whether introducing a mediator into a substructure increases or decreases the stability of the overall structure. We prove theorems that quantify the stability of a political structure when $n$ mediators are introduced, either one by one or simultaneously. We also examine how the stability is affected when a single agent is split into two. In addition, stability is expressed in terms of the $h$-vector, and special attention is given to a class of political structures modeled by shellable simplicial complexes. In the latter context, we analyze weighted political structures and examples of political structures modeled by independence complexes of graphs. This approach provides a rigorous, stepwise analysis of stability under different structural modifications, showing how the combinatorial and topological properties of the simplicial complex govern the structure’s stability.
@@ 1. Introduction
In the specific case that is relevant to us here, the authors in [15] establish a natural correspondence (see the text following Definition 2.2) between the political structure (Definition 2.1) on the one hand and the simplicial complex (Definition 2.2) on the other.
@@
In this paper, we investigate the effects of introducing a mediator into a substructure of a political structure, motivated by an open question posed after [15, Proposition 3.15], as well as the effects of introducing $n$ mediators into a political structure. We also investigate what happens when instead of merging agents we have splitting of an agent - in other words, when a vertex in a simplicial complex is split into two new vertices. Such situations are natural to observe, bearing in mind, for example, the splitting of political parties into new parties. Furthermore, we describe the stability of a political structure modeled by a simplicial complex via the $h$-vector of the complex, with particular attention to the case of shellable complexes, which are considered primarily for their topological and combinatorial properties, but we show that they also possess substantive social-scientific relevance.
In the end, we model political structures as independence complexes of graphs, and illustrate this with several examples, showing how tools from topological combinatorics can be applied to investigate their stability.
This stepwise, combinatorial, and topological approach provides a rigorous framework to quantify stability in political structures modeled by simplicial complexes, with previous works – supplemented here by new results – demonstrating the relevance and applicability of this methodology.
@@ 2. Overview of some basic concepts and results
The cited literature - especially [1, 6, 15] - provides a detailed introduction to the terminology of political structures, as well as the necessary background on simplicial complexes. Accordingly, here we will limit ourselves to a brief review of key concepts and results that are directly relevant to the problems under consideration, and construct several illustrative examples.
@@
Any simplex in the complex is also called a face of the complex.
@@
Given a $d$-dimensional political structure $P$ with the set of agents $V=\{v_{0},v_{1},\ldots,v_{k}\}$, where $k\geq 1$,
@@
    where $\operatorname{st}(v_{i})$ is the star of $v_{i}$, i.e. the collection of all simplices of $P$ containing $v_{i}$, and $|\cdot|$ denotes the cardinality.
For $k=0$, we define by convention $\operatorname{stab}(P)=0$ and $\operatorname{via}(v_{0})=0$. In the sequel, we assume that $P$ consists of at least two agents.
@@
We note that the intuitive concept of stability of a political structure cannot be fully captured by a single numerical index; however, it is natural in this context to adopt the notion introduced in the works upon which we build. We have addressed this issue in several details in the final section.
@@
We provide the following example to illustrate and motivate several concepts discussed later in the text. We will refer to it several times.
Example 2.8.
Let $P_{a,b;s}$ be a simplicial complex modelling a political structure, obtained by gluing an (a1)(a-1)-dimensional simplex $\Delta^{a-1}$ and an (b1)(b-1)-dimensional simplex $\Delta^{b-1}$ along a common (s1)(s-1)-dimensional simplex $\Delta^{s-1}$, where $1\leq s<a\leq b$. In short,
$P_{a,b;s}=\Delta^{a-1}\bigcup_{\Delta^{s-1}}\Delta^{b-1}.$
Figure 3 depicts $P_{3,4;2}$. A direct computation yields
$\operatorname{stab}(P_{a,b;s})=\frac{2^{a}-a+2^{b}-b-2^{s}+s-1}{2^{a+b-s}-a-b+s-1}.$
Although $P_{a,b;s}$ has a large number of simplices when $a$ or $b$ is large, its stability can be computed explicitly. Moreover, $\operatorname{stab}(P_{a,b;s})$ is an increasing function of $s$ for fixed $a$ and $b$.
The viability of a vertex in $P_{a,b;s}$ takes just two (for $a=b$) or three (for $a\neq b$) values, namely
$\frac{2^{a-1}-1}{2^{a+b-s-1}-1},\frac{2^{b-1}-1}{2^{a+b-s-1}-1},\frac{2^{a-1}+2^{b-1}-2^{s-1}-1}{2^{a+b-s-1}-1}.$
Figure 3. The political structure $P_{3,4;2}$
@@
The shifted $f$-polynomial of the structure $P_{a,b;s}$ from Example 2.8 is
$F_{P_{a,b;s}}(x)=\sum_{i=2}^{b}\left(\binom{a}{i}+\binom{b}{i}-\binom{s}{i}\right)x^{i-1},$
where, by convention, $\binom{n}{k}=0$ for $k>n$.
@@ 3. Effects of introducing a mediator into the substructure of a political structure
We made a simple observation that identifies a condition under which the stability of the structure increases or decreases, not restricted to the case $G\geq k+1$, thereby providing a more complete answer to the aforementioned question from [15]. We state it as an immediate corollary to the previous proposition.
@@
Therefore, $G$ is a characteristic indicating how benevolent (acceptable) a mediator must be in order to increase stability. Like the main notion $\operatorname{stab}(P)$, it depends only on the total number of vertices and the number of simplices of $P$. In Remark 3.4, we give a consequence and an interpretation of this result, which actually provide a complete answer to the question from [15] that we consider.
@@
where $S\in\{k+1,k+2,\ldots,2^{k+1}-1\}$. The function $g(S)$ is strictly decreasing. Consequently, as $S$ increases, the number of unaffected simplices $G<g(S)$ for which the introduction of a mediator leads to an increase in system stability decreases. In other words, the more “developed” the structure – i.e., the greater the number of simplices or pre-existing compatibilities among agents – the larger the subcomplex (in terms of cardinality) into which the mediator must be introduced in order to achieve an increase in overall structure stability.
@@
Remark 3.4.
Given a political structure $P$ with $k+1$ vertices,
$\operatorname{stab}(P)\in\left\{0,\frac{1}{2^{k+1}-k-2},\frac{2}{2^{k+1}-k-2},\ldots,1-\frac{1}{2^{k+1}-k-2},1\right\}\subset[0,1].$
These points are equally spaced within the unit interval $[0,1]$. Let $G$ denote the number of unaffected simplices from Proposition 3.1 and Corollary 3.2. The case $G\geq k+1$ is clear. Let $G\in\{0,1,\ldots,k+1\}$. Then the points $1-\frac{G}{k+1}$ partition the interval $[0,1]$ into $k+1$ congruent subintervals. It is easy to see that the condition
$G>\frac{(k+1)(2^{k+1}-S-1)}{2^{k+1}-k-2}$
from Corollary 3.2 is equivalent to
$1-\frac{G}{k+1}<\operatorname{stab}(P).$
In other words, the stability increases after introducing a mediator if and only if
$\operatorname{stab}(P)<1-\frac{G}{k+1},$
that is, if and only if the point representing the stability of $P$ lies to the left of the point representing $1-\frac{G}{k+1}$. The right end $G=k+1$ corresponds to the result from Proposition 3.1, while the case $G=0$ is trivial. We have already shown in Example 3.3-c) that the points $\operatorname{stab}(P)$ and $1-\frac{G}{k+1}$ never coincide.
We also consider the following interpretation of the previous results to be useful. Our goal here is to introduce a mediator in such a way that the stability of the political structure increases; in other words, we introduce a cone over a subcomplex so that the stability of the resulting complex is greater than that of the initial complex. We seek the smallest possible subcomplex. Only a small number of simplices may remain unaffected, and how many do so depends on the number of vertices and the stability of the initial complex. The lower the stability of the initial complex, the more simplices may remain unaffected.
@@ 4. Effects of introducing many mediators into a political structure
As previously described, introducing a single mediator $v$ into a political structure $P$ yields the cone $CP=P*\{v\}$ on $P$. Introducing a mediator increases the stability of the structure, that is, $\operatorname{stab}(P)\leq\operatorname{stab}(CP)$. Introducing simultaneously two mediators $v$ and $w$ into $P$ yields the suspension $\Sigma P=P*\{v,w\}$, where $\{v,w\}$ is just the set of vertices $v$ and $w$. The authors in [15] claim that $\operatorname{stab}(P)\leq\operatorname{stab}(\Sigma P)$; however, this need not be true. Their argument rests on the fact that the suspension can be seen as the pushout of two copies of the cone $CP$ along $P$, together with the claim that a pushout increases stability. But the pushout does not in general increase stability. It does so only when no new vertices are introduced, i.e. when we work with $P_{1}\coprod_{S}P_{2}$, where $S$ is a subcomplex of both $P_{1}$ and $P_{2}$, and $P_{2}$ introduces no new vertices ($V(P_{2})\subseteq V(P_{1})$). In the case of the suspension, of course, new vertices are necessarily added. In general, one often encounters
@@
These two cases, in fact, represent the extreme ways of introducing $n$ mediators. In principle, one could mix the two approaches, introducing some mediators sequentially and others simultaneously; such a mixed case can then be obtained by successive application of these two extreme cases.
@@
Example 4.1.
The structure $P_{a,b;s}$ from Example 2.8 can be obtained by introducing $s$ mediators one by one into the disjoint union
$\Delta^{a-s-1}\coprod\Delta^{b-s-1}.$
@@
The number $\frac{k+2}{2^{k+1}}$ is small for large $k$. Therefore, introducing many mediators does not significantly increase the stability of the structure, especially in the case of structures with a larger number of agents. In this context, one may pose the following question, which would require a more formal framework to answer: given that introducing mediators brings with it a cost while increasing stability yields a benefit, how many mediators should be introduced for their introduction to be justified?
@@ 5. Splitting and stability
Assume that $v$ is an agent (a vertex) of $P$ and that $v$ splits into two agents $a$ and $b$ (new vertices). Each of these new agents forms new coalitions, which can be described as the links $L_{a}\subset P$ and $L_{b}\subset P$ of the vertices $a$ and $b$, respectively. Therefore, the resulting political structure is
$P_{v}^{\prime}=(P\setminus\{v\})\amalg\left(a*L_{a}\right)\amalg\left(b*L_{b}\right).$
@@
Therefore, $X$ is a characteristic described using the notions introduced in [15]. In addition, in this case one could proceed similarly to the discussion in Remark 3.4.
@@
Example 5.4.
The structure $P_{a,b;s}$ from Example 2.8 can be obtained from $P_{a,b;s+1}$ by splitting one of the common agents in such a way that one of the new agents is compatible with all faces of $\Delta^{a-1}$, while the other is compatible with all faces of $\Delta^{b-1}$. Since the function $\operatorname{stab}(P_{a,b;s})$ is increasing with respect to $s$, we have
$\operatorname{stab}(P_{a,b;s})<\operatorname{stab}(P_{a,b;s+1}).$
@@ 6. Shellable political structures and their stability
The structure $P_{a,b;s}$ from Example 2.8, and in particular the fact that the maximal stability of $P_{b,b;s}$ is attained when s=b1s=b-1, is one of the motivations for introducing the concept of shellability in our context.
@@
As an example, the boundary of every simplicial polytope is shellable. From a topological perspective, a shellable simplicial complex of dimension $d$ is homotopy equivalent to a wedge of $d$-dimensional spheres or is contractible.
Remark 6.3.
In summary, a shellable political structure $P$ of dimension $d$ satisfies the following two conditions:
  • all maximal viable configurations of $P$ have the same number of agents, namely $d+1$;
  • for any two maximal viable configurations $F$ and $F^{\prime}$ of $P$, there exists a path $F=F_{1},F_{2},\ldots,F_{n}=F^{\prime}$ such that $|F_{i}\cap F_{i+1}|=d$ for each $i$.
This can be interpreted as follows. Any maximal viable configuration (coalition) can be obtained from any other through small changes, where one agent is replaced by another. At each step, the set remains a maximal viable configuration, allowing one to move from one coalition to another via small, incremental steps without drastic disruptions.
@@
Example 6.4.
The $h$-vector of the simplicial complex $P_{b,b-1;b-2}$, which is not shellable, is $(1,1,-1,0,\ldots,0)$.
The simplicial complex $P_{b,b;b-1}$ is shellable. Its $h$-vector is $(1,1,0,\ldots,0)$.
If the political structure under consideration is shellable, we observe that formula (4) becomes particularly useful. In that case, we can easily track how the stability changes by adding or deleting some facets (keeping the shellability). Adding (or removing) a new facet of type $i>1$ (there are no new vertices, i.e. agents), increases (or decreases) the corresponding entry $h_{i}$ by one. Consequently, the stability of $P$ increases (or decreases) by
@@
Consider what happens if we glue a facet that brings one new vertex, i.e., a facet of type $1$. Assume that
$P=F_{1}\cup F_{2}\cup\cdots\cup F_{j}$
and let $F_{j+1}$ be a $d$-dimensional simplex of type 1, where $P$ has $k+1$ agents. If $\widetilde{S}$ is the total number of positive-dimensional simplices in $P$, then $\operatorname{stab}(P)=\frac{\widetilde{S}}{2^{k+1}-k-2}$. If the dimension of $P$ is $d$, then for
@@
If all complexes $P_{j}$ are shellable, and if $P_{j+1}$ is obtained from the previous one by adding a few new faces, we can easily calculate the stabilities and the weighted stability. This in itself is a reasonable motivation to consider shellable structures in this context.
@@ 7. The independence complex of a graph as a political structure
Here we describe a natural way to model political structures using graphs and provide two illustrative examples.
Consider a graph $G=(V,E)$ with $k+1$ vertices. Assume that the vertices of $G$ represent the agents. Let any two agents connected by an edge in $G$ be in conflict and therefore unable to appear together in any viable configuration. In this setting, the set of all viable configurations corresponds precisely to the independence complex $I(G)$ of the graph $G$, which is the simplicial complex whose faces are the independent sets of vertices of $G$.
For a formal definition and more on independence complexes, including their connection to shellability, see [9].
@@
Consider $k+1=2n$ agents labeled by the set $\{1,2,\ldots,n\}\cup\{-1,-2,\ldots,-n\}$. Let each agent $i$ be in conflict solely with agent i-i. These conflicts can then be represented by a graph consisting of $n$ edges, each connecting $i$ and i-i for all $i\in[n]$. The corresponding political structure is $P=\partial C^{\triangle}_{n-1}$, the boundary of (n1)(n-1)-dimensional crosspolytope $C^{\triangle}_{n-1}$ (the dual of (n1)(n-1)-cube). The complex $\partial C^{\triangle}_{n-1}$ has $2^{n}$ facets, each corresponding to an ordered partition of $[n]$ into two subsets. For a partition $n]=A\sqcup B$, the facet associated to $(A,B)$ is defined by
@@ 8. Further questions
    According to the definition of stability of a political structure, two structures with the same number of vertices and the same total number of $(\geq 1)$-dimensional simplices will have the same stability. However, these structures need not be topologically equivalent. Therefore, the stability defined in this manner is not a topological invariant, which makes it problematic if one aims to approach the problem from a more topological perspective.
    Structures with $k+1$ vertices are partitioned into $2^{k+1}$ “stability classes”. Even the dimension of the structure does not appear as something relevant.
@@
    Another idea to distinguish structures, at least by their $f$-vectors, is to define
    $\widehat{\operatorname{stab}}(P)=\frac{1}{f_{0}}\sum_{i=0}^{\dim{P}}\frac{f_{i}}{\binom{f_{0}}{i+1}}$.
    With this definition, for a fixed total number of simplices, the function $\widehat{\operatorname{stab}}$ favors structures that have more simplices of lower dimension, in the sense that they attain a higher value of $\widehat{\operatorname{stab}}(P)$.